OpenStudy (anonymous):

Will someone explain this to me? a. The vertex form of the equation of a vertical parabola is given by , where (h, k) is the vertex of the parabola and the absolute value of p is the distance from the vertex to the focus, which is also the distance from the vertex to the directrix.

OpenStudy (anonymous):

You will use the GeoGebra geometry tool to create a vertical parabola and write the vertex form of its equation. i:i. Mark the focus of the parabola you are going to create at F(6, 4). Draw a horizontal line that is 6 units below the focus. This line will be the directrix of your parabola. What is the equation of the line? i am super confused on what to do

OpenStudy (anonymous):

hey do u still need help?

OpenStudy (anonymous):

"focus of the parabola at (6, 4). horizontal line that is 6 units below the focus will be the directrix" directrix: y = -2 Determine the orientation of the parabola: The directrix is horizontal, so the parabola is vertical. The focus lies above the directrix, so the parabola opens upwards. Equation of up-opening parabola:  y = a(x - h)² + k The vertex is halfway between focus and directrix. vertex (6, 1) h = 6 k = 1 p = distance between focus and vertex = 3 a = 1/(4p) = 1/12 The equation of the parabola: y = (1/12)(x - 6)² + 1

OpenStudy (anonymous):

thank you